OpenStudy (anonymous):

Stable High School has a total of 112 boys and girls who play sports. If the number of boys, b is 16 more than twice the number of girls, g, how many boys play sports at this high school? Which system of equations would help you solve this problem? b + g = 112 b = 2g + 16 b + g = 112 g = 2b + 16 b − g = 112 b = 2g − 16 None of the systems above would solve this problem.

OpenStudy (lilg132):

72

OpenStudy (anonymous):

I did this like 5 time to make sure im right. i came up with the first one everytime but i still think its wrong

OpenStudy (lilg132):

(b + g) / 2 + 16 = 72

OpenStudy (lilg132):

b = 2g + 16 yes thats correct

OpenStudy (lilg132):

b + 16 = 2g is wrong

OpenStudy (phi):

First, yes you did it correctly. The statment number of boys, b is 16 more than twice the number of girls, g, translates as follows: b is becomes b = (is translates to equal most of the time in these problems) 16 more than becomes 16+ (and 16 less than would be -16 +) so we have b=16+ so far twice the number of girls g becomes 2 times the number of girls, or 2g so we finally get b= 16+2g And the only choice with this equation is the first choice

OpenStudy (anonymous):

thank you both very much!

OpenStudy (anonymous):

ok